OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=4/5, (4,-2) y=

OpenStudy (anonymous):

m=4/5, (4,-2) y = mx +c [substitute known values into general equation] -2 = 4/5(4) + c -2 = 16/5 + c Subtract 16/5 from both sides to reveal c -36/5 = c Now you can write the equation of the given line y = 4x/5-36/5